OpenStudy (deadshot):

How do I solve Solve 8^(x - 2) = 3?

OpenStudy (deadshot):

\[8^{x-2}=3\]

OpenStudy (a_clan):

Use logarithms

OpenStudy (unklerhaukus):

take log to base eight of both sides

OpenStudy (deadshot):

so \[\log _{8}x-2=\log _{8}3\]?

OpenStudy (unklerhaukus):

almost

OpenStudy (unklerhaukus):

\[8^{x-2}=3\\\log _{8}8^{x-2}=\log _{8}3\\%{x-2}=\log _{8}3\]

OpenStudy (anonymous):

yes. you do this because, the variable we are solving for is in the exponent and we have to bring it down. The opposite of exponentiation is logarithm!